# A Survey on Large Language Models for Code Generation

### v1

@@ 4.10.3. LLM-as-a-Judge
Table 8. The overview of code assistant applications powered by large language models (LLMs). The column labeled ‘PLs’ and ‘IDEs’ indicate programming languages and integrated development environments, respectively (Zan et al., 2023).
@@ 5. Challenges & Opportunities
According to our investigations, the LLMs have revolutionized the paradigm of code generation and achieved remarkable performance. Despite this promising progress, there are still numerous challenges that need to be addressed. These challenges are mainly caused by the gap between academia and practical development. For example, in academia, the HumanEval benchmark has been established as a de facto standard for evaluating the coding proficiency of LLMs. However, many works have illustrated the evaluation of HumanEval can’t reflect the scenario of practical development (Jimenez et al., 2023; Du et al., 2024; Liu et al., 2024b; Ding et al., 2024). In contrast, these serious challenges offer substantial opportunities for further research and applications. In this section, we pinpoint critical challenges and identify promising opportunities, aiming to bridge the research-practicality divide.
Enhancing complex code generation at repository and software scale. In practical development scenarios, it often involves a large number of complex programming problems of varying difficulty levels. While LLMs have shown proficiency in generating function-level code snippets, these models often struggle with more complex, unseen programming problems, repository- and software-level problems that are commonplace in real-world software development. To this end, it requires strong problem-solving skills in LLM beyond simply functional-level code generation. For example, AlphaCode (Li et al., 2022a) achieved an average ranking in the top 54.3% in programming competitions where an understanding of algorithms and complex natural language is required to solve competitive programming problems. (Jimenez et al., 2023) argues that existing LLMs can’t resolve real-world GitHub issues well since the best-performing model, Claude 2, is able to solve a mere 1.96% of the issues. The reason for poor performance is mainly attributed to the weak reasoning capabilities (Huang and Chang, 2022), complex internal- and external- dependencies (Bairi et al., 2023), and context length limitation of LLMs (Bairi et al., 2023). Therefore, the pursuit of models that can handle more complex, repository- and software-level code generation opens up new avenues for automation in software development and makes programming more productive and accessible.
@@
Curating high-quality code data for pre-training and fine-tuning of LLMs. The efficacy of LLMs largely depends on the quality and diversity of code datasets used during pre-training and fine-tuning phases (Zhou et al., 2024; Köpf et al., 2024; Wettig et al., 2024). Currently, there is a scarcity of large, high-quality datasets that encompass a wide range of programming tasks, styles, and languages. This limitation constrains the ability of LLMs to generalize across unseen programming tasks, different coding environments, and real-world software development scenarios. The development of more sophisticated data acquisition techniques, such as automated code repositories mining (Linstead et al., 2007), advanced filtering algorithms, and code data synthesis (Liu et al., 2024a) (see Section 4.2), can lead to the creation of richer datasets. Collaborations with industry partners (e.g., GitHub) could also facilitate access to proprietary codebases, thereby enhancing the practical relevance of the training material. Furthermore, the adoption of open-source models for dataset sharing can accelerate the collective effort to improve the breadth and depth of code data available for LLM research.
@@
Ensuring code safety and aligning LLM outputs with human coding preferences. Ensuring the safety and security of code generated by LLMs is a paramount concern, as is their ability to align with human preferences and ethical standards. Current models may inadvertently introduce vulnerabilities or generate code that does not adhere to desired norms (Chen et al., 2021; Yang et al., 2024). Research into the integration of formal verification tools within the LLM pipeline can enhance the safety of the produced code. Additionally, developing frameworks for alignment learning that capture and reflect human ethical preferences can ensure that the code generation process aligns with societal values (Ouyang et al., 2022; Qi et al., 2023). Transparent and explainable AI methodologies can also contribute to building trust in the LLM-generated code by making the decision-making process more accessible to developers.
@@ 6. Conclusion
In this survey, we provide a systematic literature review, serving as a valuable reference for researchers investigating the cutting-edge progress in LLMs for code generation. A thorough introduction and analysis for data curation, the latest advances, performance evaluation, and real-world applications are illustrated. In addition, we present a historical overview of the evolution of LLMs for code generation in recent years and offer an empirical comparison using the widely recognized HumanEval and MBPP benchmarks to highlight the progressive enhancements in LLM capabilities for code generation. Critical challenges and promising opportunities regarding the gap between academia and practical development are also identified for future investigation. Furthermore, we have established a dedicated resource website to continuously document and disseminate the most recent advances in the field. We hope this survey can contribute to a comprehensive and systematic overview of LLM for code generation and promote its thriving evolution. We optimistically believe that LLM will ultimately change all aspects of coding and automatically write safe, helpful, accurate, trustworthy, and controllable code, like professional programmers, and even solve coding problems that currently cannot be solved by humans.

### v2

@@ 5.11. Code LLMs Alignment
Table 12. The overview of code assistant applications powered by LLMs. The column labeled ‘PLs’ and ‘IDEs’ indicate programming languages and integrated development environments, respectively (Zan et al., 2023).
@@
Figure 19. An exemplar of GitHub Copilot to demonstrate how to use development tools powered by LLMs, including powerful GPT 4o, o1-preview (Preview), and o1-mini (Preview). To illustrate its capabilities, we input the description of the “5. Longest Palindromic Substring” problem from LeetCode into Copilot’s chat box. The code generated by Copilot is then submitted to the online judge platform, where it is successfully accepted.
@@ 5.12. Applications
To illustrate the use of development tools powered by LLMs, we employ GitHub Copilot within Visual Studio Code (VS Code) as our example. Note that
As depicted in the upper section of Figure 19, users can interact with Copilot through the chat box in the lower left corner, where they can inquire about various coding-related tasks. This feature is now supported by the advanced capabilities of GPT-4o, o1-preview (Preview), and o1-mini (Preview). From the generated content, Copilot demonstrates the ability to plan solutions to coding problems. It can write code and subsequently explain the generated code to enhance user comprehension. Within the right-side workspace, users can engage in inline chat conversations to generate or refactor source code, conduct code explanations, fix coding errors, resolve issues encountered during terminal command executions, produce documentation comments, and generate unit tests. To illustrate its capabilities, we input the description of the “5. Longest Palindromic Substring” problem from LeetCode into Copilot’s chat box. The code generated by Copilot is then submitted to the online judge platform, where it is successfully accepted, as shown at the lower section of Figure 19.
@@ 6. Challenges & Opportunities
According to our investigations, the LLMs have revolutionized the paradigm of code generation and achieved remarkable performance. Despite this promising progress, there are still numerous challenges that need to be addressed. These challenges are mainly caused by the gap between academia and practical development. For example, in academia, the HumanEval benchmark has been established as a de facto standard for evaluating the coding proficiency of LLMs. However, many works have illustrated the evaluation of HumanEval can’t reflect the scenario of practical development (Jimenez et al., 2023; Du et al., 2024; Liu et al., 2024c; Ding et al., 2024). In contrast, these serious challenges offer substantial opportunities for further research and applications. In this section, we pinpoint critical challenges and identify promising opportunities, aiming to bridge the research-practicality divide.
Enhancing complex code generation at repository and software scale. In practical development scenarios, it often involves a large number of complex programming problems of varying difficulty levels (Zhang et al., 2022; Li et al., 2022a). While LLMs have shown proficiency in generating function-level code snippets, these models often struggle with more complex, unseen programming problems, repository- and software-level problems that are commonplace in real-world software development. To this end, it requires strong problem-solving skills in LLM beyond simply functional-level code generation. For example, AlphaCode (Li et al., 2022a) achieved an average ranking in the top 54.3% in programming competitions where an understanding of algorithms and complex natural language is required to solve competitive programming problems. (Jimenez et al., 2023) argues that existing LLMs can’t resolve real-world GitHub issues well since the best-performing model, Claude 2, is able to solve a mere 1.96% of the issues. The reason for poor performance is mainly attributed to the weak reasoning capabilities (Huang and Chang, 2022), complex internal- and external- dependencies (Bairi et al., 2023), and context length limitation of LLMs (Bairi et al., 2023). Therefore, the pursuit of models that can handle more complex, repository- and software-level code generation opens up new avenues for automation in software development and makes programming more productive and accessible.
@@
Curating high-quality code data for pre-training and fine-tuning of LLMs. The efficacy of LLMs largely depends on the quality and diversity of code datasets used during pre-training and fine-tuning phases (Zhou et al., 2024; Köpf et al., 2024; Wettig et al., 2024). Currently, there is a scarcity of large, high-quality datasets that encompass a wide range of programming tasks, styles, and languages. This limitation constrains the ability of LLMs to generalize across unseen programming tasks, different coding environments, and real-world software development scenarios. The development of more sophisticated data acquisition techniques, such as automated code repositories mining (Linstead et al., 2007), advanced filtering algorithms, and code data synthesis (Liu et al., 2024b) (see Section 5.2), can lead to the creation of richer datasets. Collaborations with industry partners (e.g., GitHub) could also facilitate access to proprietary codebases, thereby enhancing the practical relevance of the training material. Furthermore, the adoption of open-source models for dataset sharing can accelerate the collective effort to improve the breadth and depth of code data available for LLM research.
@@
Ensuring code safety and aligning LLM outputs with human coding preferences. Ensuring the safety and security of code generated by LLMs is a paramount concern, as is their ability to align with human preferences and ethical standards. Current models may inadvertently introduce vulnerabilities or generate code that does not adhere to desired norms (Chen et al., 2021; Yang et al., 2024a). Research into the integration of formal verification tools within the LLM pipeline can enhance the safety of the produced code. Additionally, developing frameworks for alignment learning that capture and reflect human ethical preferences can ensure that the code generation process aligns with societal values (Ouyang et al., 2022; Qi et al., 2023). Transparent and explainable AI methodologies can also contribute to building trust in the LLM-generated code by making the decision-making process more accessible to developers.
@@ 7. Conclusion
In this survey, we provide a systematic literature review, serving as a valuable reference for researchers investigating the cutting-edge progress in LLMs for code generation. A thorough introduction and analysis for data curation, the latest advances, performance evaluation, ethical implications, environmental impact, and real-world applications are illustrated. In addition, we present a historical overview of the evolution of LLMs for code generation in recent years and offer an empirical comparison using the widely recognized HumanEval, MBPP, and the more practical and challenging BigCodeBench benchmarks to highlight the progressive enhancements in LLM capabilities for code generation. Critical challenges and promising opportunities regarding the gap between academia and practical development are also identified for future investigation. Furthermore, we have established a dedicated resource website to continuously document and disseminate the most recent advances in the field. We hope this survey can contribute to a comprehensive and systematic overview of LLM for code generation and promote its thriving evolution. We optimistically believe that LLM will ultimately change all aspects of coding and automatically write safe, helpful, accurate, trustworthy, and controllable code, like professional programmers, and even solve coding problems that currently cannot be solved by humans.